[figure: 9.7 cm]

PREPARATION OF HIGH QUALITY MBE-GROWN Au/Mo/Al_{2}O_{3}(11-20) THIN FILMS FOR EPITAXY OF ULTRATHIN COBALT LAYERS

L.T. Baczewski^{a}, A. Wawro^{a}, J. Swieklo^{b}, A. Maziewski^{b} aInstitute of Physics, Polish Academy of Sciences, Warsaw. Poland bInstitute of Physics, University of Bialystok, Poland

Magnetic properties of ultrathin metallic layers depend on their surface quality which is closely related to a roughness of buffer layer. In this work we have studied the quality of molecular beam epitaxy (MBE) grown Au thin films deposited on a sapphire substrate Al_{2}O_{3} oriented in [11-20] direction covered by Mo buffer layer and its dependence on deposition conditions and annealing temperature. The structure of the film was checked in-situ with reflection high-energy electron diffraction (RHEED) and Auger spectroscopy and ex-situ with scanning tunneling microscope (STM) and atomic force microscope (AFM). Procedure for improvement of a sample surface quality was developed for exsitu STM/AFM study. The dependencies of average grain size and terraces width on annealing temperature were obtained for Mo buffer layer and Au thin film grown on it. The aim of this work is to obtain optimal growth and preparation conditions for epitaxy of ultra-thin Co films with perpendicular magnetic anisotropy grown on a gold buffer layer. This work was partially supported by the Polish Committee for Scientific Research under grants no. 8T11B05309 and 2P03B06515 and ESF NANOMAG Program.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : L.T. Baczewski

Address for correspondence : Institute of Physics, Polish Academy of Sciences, Al. Lotnikow 32/46, 02-668 Warsaw, Poland

Email address : bacze@ifpan.edu.pl